how to find the "y"..??
x=-6
y=-2/3x+1

Answer by funmath(2933)

how to find the "y"..??
E1: x=-6
E2: y=-2/3x+1
:
Let x in E2 be -6 and solve for y.
y=(-2/3)(-6)+1
y=12/3+1
y=4+1
y=5
:
Happy Calculating!!!